18. suppose that the random variable X has a continuous uniform distribution over [ 10, 20 ]. Find P (15 is less than or equal to X is less than or equal to 19 )

Possible answers :
A 0.30, B 0.35, C 0.40, D 0.45
Answer #1

here a = 10 and b = 20
P(X < x) = (x - a)/(b -a)

P(15 <= X <= 19)
= P(X <= 19) - P(X <= 15)
= (19 - 10)/(20 - 10) - (15 - 10)/(20 - 10)
= 0.9 - 0.5
= 0.4

option C